# Power to the Clients: Federated Learning in a Dictatorship Setting

### v1

@@ 4.2 Collaborative Dictator Clients
In this section, we extend the single dictator client scenario to a group of $P$ dictator clients acting in coordination. As illustrated in Figure 1(c), these clients collaborate to suppress the influence of all others while preserving their own contributions, relying only on inter-client communication. As discussed in Appendix C, they do not require prior knowledge of the server’s learning rate—it can be accurately estimated after a single training round. Let $\mathcal{P}\subset\mathcal{N}$ denote the set of $P$ collaborating dictator clients, where $1<P<N$, capable of communicating with each other. These clients coordinate their updates according to Algorithm 2 so that the global model evolves as if only they had participated in training. Each client in $\mathcal{P}$ maintains a synchronized local model, denoted as $\hat{\theta}^{\mathcal{P}}_{t}$, representing the model state at iteration $t$ under their exclusive contributions from the start. At each round, every dictator client $k\in\mathcal{P}$ submits the following crafted update to the server, effectively nullifying the impact of the remaining NPN-P clients in $\mathcal{N}\setminus\mathcal{P}$:
$M^{k}_{t}=\nabla{\mathcal{L}_{k}(\hat{\theta}^{\mathcal{P}}_{t})}-\left(\frac{\hat{\theta}^{\mathcal{P}}_{t-1}-\theta_{t}}{P\eta}-\nabla{\mathcal{L}_{k}(\hat{\theta}^{\mathcal{P}}_{t-1}})\right),\forall k\in\mathcal{P}.$
Afterwards, the clients exchange gradients to jointly compute the next local model state, $\hat{\theta}^{\mathcal{P}}_{t+1}$. As long as all $P$ clients in $\mathcal{P}$ follow this protocol and continues to share gradients for updating the collective local model, the global model will converge as if only the clients in $\mathcal{P}$ had trained it. A formal proof of this outcome is provided in Appendix D. Next, we turn our attention to more intricate interactions that emerge in FL systems with the presence of such dictator clients.
@@
5 Competition and Collusion Among Dictator Clients
In this section, we explore the nuanced interactions that can arise among dictator clients in FL systems. We begin by examining a competitive setting where every participating client independently aims to become the sole dictator and dominate the global model—an extreme yet insightful scenario that is discussed in Section 5.1. Furthermore, in Section 5.2, we explore a more collaborative dynamic, where multiple dictator clients form alliances. We investigate whether such cooperation is inherently stable or if, ultimately, some clients can strategically betray their collaborators to gain a greater influence over the model.
5.1 Mutual Domination: When Every Client Seeks Control
Here, we explore the scenario where all clients independently act as dictators, each attempting to retain only its own contribution while nullifying the influence of others. In other words, each client follows the update strategy outlined in Algorithm 1. In practice, such behavior leads to a catastrophic failure of learning, with the global model failing to converge and the loss growing exponentially. We analyze the underlying reason behind this phenomenon in what follows. At iteration 0, the server sends the initialized weights $\theta_{0}$ to all clients. Each client then computes its local gradient, and the server aggregates these to update the global model as $\theta_{1}=\theta_{0}-\eta\sum\nolimits_{n=1}^{N}\nabla\mathcal{L}_{n}(\theta_{0})$. In the next iteration, the server broadcasts $\theta_{1}$ to all clients. Now, each client attempts to simulate what the global model would have been if it alone had contributed to the update. For each client $n\in\mathcal{N}$, we define $\hat{\theta}^{n}_{1}$ as the hypothetical global model after iteration 0 only if client $n$ had participated. Using this, each client computes its malicious update $M^{n}_{1}$, as defined in Section 4.1 as $M^{n}_{1}=\nabla{\mathcal{L}_{n}}(\hat{\theta}^{n}_{1})-\left(\frac{\theta_{0}-\theta_{1}}{\eta}-\nabla\mathcal{L}_{n}(\theta_{0})\right)$. Now, we analyze the updated global model $\theta_{2}$ after the server aggregates the updates from all clients in the second iteration:
$\theta_{2}$ $=\theta_{1}-\eta\sum\nolimits_{n=1}^{N}M^{n}_{1}=\theta_{1}-\eta\left(\sum\nolimits_{n=1}^{N}{\nabla{\mathcal{L}_{n}}(\hat{\theta}^{n}_{1})-(\frac{\theta_{0}-\theta_{1}}{\eta}-\nabla\mathcal{L}_{n}(\theta_{0}))}\right)$
$=\theta_{1}-\eta\left(\sum\nolimits_{n=1}^{N}\nabla{\mathcal{L}_{n}}(\hat{\theta}^{n}_{1})-\frac{N(\theta_{0}-\theta_{1})}{\eta}+\sum\nolimits_{n=1}^{N}\nabla{\mathcal{L}_{n}(\theta_{0})}\right)$
$=\theta_{0}-\eta\sum\nolimits_{n=1}^{N}\nabla{\mathcal{L}_{n}(\theta_{0})}+\eta(N-1)\sum\nolimits_{n=1}^{N}\nabla{\mathcal{L}_{n}(\theta_{0})}-\eta\sum\nolimits_{n=1}^{N}\nabla{\mathcal{L}_{n}(\hat{\theta}^{n}_{1}})$
Since N2>0N-2>0 assuming that we have more than $2$ clients in the system, and the learning rate $\eta$ is a positive real number, it follows that $\eta(N-2)>0$. Consequently, from Eq. 10, it follows that when all clients act as independent dictators and send the defined malicious update, the resulting model update effectively moves in the opposite direction of intended gradient. In other words, the updating procedure resembles gradient ascent instead of gradient descent, and thereby increasing the loss rather than minimizing it. This behavior causes the model to ”unlearn” the progress made in previous iteration. Therefore, when every client behaves as an independent dictator, the global model fails to learn meaningful representations and make no effective progress. it unlearns the knowledge acquired in the previous iteration. Therefore, in the scenario where every client is an independent dictator, the global model will learn almost nothing. Our empirical results, presented in Section 6.3, confirms this breakdown in learning in practice.
5.2 Betrayal in Collaboration: Strategic Cheating Among Dictator Clients
Here, we show that even collaborative dictators—those collaborating to erase other participants’ contributions—may ultimately betray one another. For simplicity, we focus on a setting where the set of collaborative dictator clients is $\mathcal{P}=\{1,2\}$. As illustrated in Figure 2, we consider the case where dictator client $1$, decides to cheat its partner, client $2$, after a specific iteration $E$. While both clients initially cooperate using Algorithm 2 to jointly eliminate the influence of all other clients, we introduce Algorithm 3 (in Appendix B), which enables client $1$ to unilaterally eliminate client $2$’s contribution as well, effectively taking full control of the model.
Prior to iteration $E$, client $1$ shares its gradients with client $2$, contributing jointly to a local model $\hat{\theta}^{\mathcal{P}}_{t}$. However, simultaneously, client $1$ secretly maintains a private model $\hat{\theta}^{1}_{t}$, which simulates the evolution of the global model if only client $1$ participated in training. At each iteration, client $1$ computes a correction term $\Delta_{t}=\nabla{\mathcal{L}_{1}(\hat{\theta}^{m}_{t})}-(\nabla{\mathcal{L}_{1}(\hat{\theta}^{\mathcal{P}}_{t})}+\nabla{\mathcal{L}_{2}(\hat{\theta}^{\mathcal{P}}_{t})})$, which captures the discrepancy between acting alone and collaborating. These differences are accumulated into a cheating offset, denoted as $\mathrm{Cheating\_Update}$. At iteration $E$, client $1$ sends this accumulated update to the server in place of the expected collaborative update. This forces the server to jump to a state equivalent to one where if only client $1$ had participated throughout the training process—effectively eliminating the contribution of client $2$, despite their prior collaboration, as well as the benign clients’ influence. A formal proof of this strategy is provided in Appendix E; our empirical results in Section 6.4 confirm the effectiveness of this betrayal strategy in practice.
6 Experiments
In this section, we empirically evaluate the effectiveness of our proposed attack algorithms across both computer vision and natural language processing (NLP) tasks. For our main experiments, we focus on image classification using the MNIST (LeCun et al., 1998) and CIFAR10 (Krizhevsky et al., 2009) datasets with a simple convolutional neural network (CNN) as the global model. To maintain consistency with our theoretical framework, all experiments are conducted using stochastic gradient descent (SGD) as the optimizer. We simulate a FL environment with five clients, each assigned a disjoint and non-overlapping subset of the training data to create a highly not independent and identically distributed (non-IID) setting. Specifically, the training set is partitioned such that client 1 receives samples with labels 0 and 1, client 2 with labels 2 and 3, and so on, ensuring that each client maintains only two unique classes. Additional results for NLP tasks are provided in Appendix F.
6.1 Single Dictator Client
Table 1 reports the resulting classification accuracies across all clients’ datasets. We begin by evaluating the scenario in which a single client attempts to dominate the global model, following the attack strategy defined in Algorithm 1. As shown, the global model entirely fails to learn from the data of non-dictator clients, achieving a striking $0.00\%$ accuracy on their datasets. In contrast, the model maintains high accuracy on the dictator client’s local dataset, confirming that the attack successfully isolates and preserves only the dictator’s contribution. These results empirically validate the feasibility and effectiveness of the proposed single-client dictatorship attack algorithm described in Section 4.1. Furthermore, Figure 3 illustrates this effect by showing the global model’s loss on each client’s dataset under two settings: regular FL and the case where client 3 becomes dictator. In regular FL, losses decrease across all clients, whereas under dictatorship by client 3, only the loss corresponding to client 3’s dataset is minimized, while losses for all other clients worsen over time. This confirms that the dictator client successfully minimizes its own local loss while significantly impeding the global model’s ability to learn from the data of the remaining clients.
Method MNIST CIFAR-10
[0,1] [2,3] [4,5] [6,7] [8,9] [0,1] [2,3] [4,5] [6,7] [8,9]
Regular FL 96.18 79.25 66.84 88.12 66.38 39.04 12.51 31.07 23.74 52.59
Dictator client: 11 99.63 0.00 0.00 0.00 0.00 73.65 0.00 0.00 0.00 0.00
Dictator client: 22 0.00 93.92 0.00 0.00 0.00 0.00 65.19 0.00 0.00 0.00
Dictator client: 33 0.00 0.00 97.43 0.00 0.00 0.00 0.00 66.51 0.00 0.00
Dictator client: 44 0.00 0.00 0.00 98.91 0.00 0.00 0.00 0.00 73.98 0.00
Dictator client: 55 0.00 0.00 0.00 0.00 94.42 0.00 0.00 0.00 0.00 77.06
Table 1: Performance of the global model on each local dataset for MNIST and CIFAR-10 datasets and the single dictator client scenario.
Method MNIST CIFAR-10
[0,1] [2,3] [4,5] [6,7] [8,9] [0,1] [2,3] [4,5] [6,7] [8,9]
Regular FL 96.18 79.25 66.84 88.12 66.38 39.04 12.51 31.07 23.74 52.59
Table 2: Performance of the global model on each local dataset for MNIST and CIFAR-10 datasets and the collaborative dictator clients scenario.
6.2 Collaborative Dictator Clients
We next examine the impact of coordinated attacks involving multiple dictator clients. In this setting, two or three clients jointly follow the attack strategy, described in Algorithm 2, aiming to eliminate the influence of all other participants. Table 2 and Figure 4 summarize the outcomes. The results demonstrate that the collaborating dictator clients succeed in entirely erasing the influence of the benign clients, leading the global model to achieve $0.00\%$ accuracy on their data. Simultaneously, the global model maintains high accuracy on the data held by the collaborative dictators, indicating that it has effectively converged to a model tailored solely to their objectives. These findings further reinforce the practicality and scalability of our proposed attack strategy in multi-attacker scenarios. The coordinated behavior among the dictator clients allows them to dominate the training process, ensuring that the global model exclusively reflects their data distributions while ignoring the contributions of the remaining benign participants. The success of this attack highlights the vulnerability of FL even when malicious clients are in minority, provided they act in collaboration.
6.3 Mutual Domination
We now consider the extreme scenario where every client behaves as an independent dictator, each executing Algorithm 1 to preserve only its own contribution while nullifying the effects of all others. As established theoretically in Section 5.1, this adversarial configuration results in mutually destructive behavior, where no single client’s update can effectively influence the global model without being canceled out by others, resulting in a destructive equilibrium where no useful learning can occur. The empirical results, shown in Figure 6, strongly corroborate this. The global model fails to make progress on any client’s data; instead of converging, the loss increases rapidly and consistently across all datasets. This behavior aligns with the theoretical finding that the aggregated updates approximate a form of gradient ascent, undoing prior learning and leading to model divergence. This experiment underscores a key insight: when all clients prioritize their own influence at the expense of others, the entire system collapses. FL becomes ineffective, highlighting the need for defenses against not only isolated attackers but also adversarial groups.
6.4 Betrayal in Collaboration
In this experiment, we evaluate a scenario in which two clients, client $1$ and client $2$, initially act as collaborative dictators. While both begin by coordinating via Algorithm 2, client $1$ eventually deviates and follows the betrayal strategy outlined in Algorithm 3 (discussed in Section 5.2). This setup allows client $1$ to secretly prepare for a unilateral takeover of the model. As shown in Figure 6, at iteration 10—the predetermined betrayal point—the global model abruptly loses performance on client $2$’s dataset, while having even lower loss on client $1$’s data. This result confirms that client $1$ successfully erases not only the contributions of the benign clients, but also those of its former collaborator, client $2$. These findings empirically validate that a malicious client can strategically cooperate to gain trust, only to later betray its partners and assert full control over the global model. This highlights a critical vulnerability in FL; even collaborative adversaries can be exploited by more sophisticated attackers acting within their own group.
Figure 5: Loss functions for mutual domination scenario
Figure 6: Loss functions for betrayal in collaboration scenario
7 Practical Implications
8 Limitations
Despite its conceptual novelty, the current form of the attacks may be vulnerable to detection. In particular, the gradient updates produced by a malicious client—especially in the single-dictator setting—can deviate from those of honest participants and may be identified by common server-side defenses such as anomaly detection or norm-based filtering. As a result, while the attack effectively exposes a weakness in the aggregation process, it may lack the stealth required for practical deployment in real-world systems without further refinement or adaptation. Nevertheless, this work should be viewed as a starting point toward a broader line of research. It raises important questions about the boundaries between personalization, dictator behavior, and adversarial manipulation in federated systems. In addition to exposing a new class of client-driven attacks, our work also explores the dynamics of collaboration and strategic behavior among clients, including scenarios where multiple clients collude or compete to influence the global model. These insights contribute to a deeper understanding of how clients may cheat, cooperate, or exploit the system for individual gain—offering a foundation for analyzing real-world risks in decentralized learning environments. Future work could build on this by developing more nuanced attack strategies that are harder to detect, incorporating techniques such as norm-constrained updates, or optimization-based stealth objectives.
9 Conclusion
Appendix A Appendix
Appendix B Algorithm 3: Cheater Client

### v2

@@ III-B Collaborative Dictator Clients
In this section, we extend the single dictator client scenario to a group of $P$ dictator clients acting in coordination. As illustrated in Figure 3(c) (in Appendix A), these clients collaborate to suppress the influence of all others while preserving their own contributions, relying only on inter-client communication. As discussed in Appendix H, they do not require prior knowledge of the server’s learning rate—it can be accurately estimated after a single training round. Let $\mathcal{P}\subset\mathcal{N}$ denote the set of $P$ collaborating dictator clients, where $1<P<N$, capable of communicating with each other. These clients coordinate their updates according to Algorithm 2 (in Appendix C) so that the global model evolves as if only they had participated in training. Each client in $\mathcal{P}$ maintains a synchronized local model, denoted as $\hat{\theta}^{\mathcal{P}}_{t}$, representing the model state at iteration $t$ under their exclusive contributions from the start. At each round, every dictator client $k\in\mathcal{P}$ submits a crafted update to the server, effectively nullifying the impact of the remaining NPN-P clients in $\mathcal{N}\setminus\mathcal{P}$, defined as $M^{k}_{t}=\nabla{\mathcal{L}_{k}(\hat{\theta}^{\mathcal{P}}_{t})}-\left(\frac{\hat{\theta}^{\mathcal{P}}_{t-1}-\theta_{t}}{P\eta}-\nabla{\mathcal{L}_{k}(\hat{\theta}^{\mathcal{P}}_{t-1}})\right),\forall k\in\mathcal{P}$. Afterwards, the clients exchange gradients to jointly compute the next local model state, $\hat{\theta}^{\mathcal{P}}_{t+1}$. As long as all $P$ clients in $\mathcal{P}$ follow this protocol and continues to share gradients for updating the collective local model, the global model will converge as if only the clients in $\mathcal{P}$ had trained it. A formal proof of this outcome is provided in Appendix I. Next, we turn our attention to more intricate interactions that emerge in FL systems with the presence of such dictator clients.
IV Experiments
In this section, we empirically evaluate the effectiveness of our proposed attack algorithms across both computer vision and natural language processing (NLP) tasks. For our main experiments, we focus on image classification using the MNIST [9] and CIFAR10 [7] datasets with a simple convolutional neural network (CNN) as the global model. To maintain consistency with our theoretical framework, all experiments are conducted using stochastic gradient descent (SGD) as the optimizer. We simulate a FL environment with five clients, each assigned a disjoint and non-overlapping subset of the training data to create a highly not independent and identically distributed (non-IID) setting. Specifically, the training set is partitioned such that client 1 receives samples with labels 0 and 1, client 2 with labels 2 and 3, and so on, ensuring that each client maintains only two unique classes. Additional results for NLP tasks are provided in Appendix K.
IV-A Single Dictator Client
Table I reports the resulting classification accuracies across all clients’ datasets. We begin by evaluating the scenario in which a single client attempts to dominate the global model, following the attack strategy defined in Algorithm 1. As shown, the global model entirely fails to learn from the data of non-dictator clients, achieving a striking $0.00\%$ accuracy on their datasets. In contrast, the model maintains high accuracy on the dictator client’s local dataset, confirming that the attack successfully isolates and preserves only the dictator’s contribution. These results empirically validate the feasibility and effectiveness of the proposed single-client dictatorship attack algorithm described in Section III-A. Furthermore, Figure 1 illustrates this effect by showing the global model’s loss on each client’s dataset under two settings: regular FL and the case where client 3 becomes dictator. In regular FL, losses decrease across all clients, whereas under dictatorship by client 3, only the loss corresponding to client 3’s dataset is minimized, while losses for all other clients worsen over time. This confirms that the dictator client successfully minimizes its own local loss while significantly impeding the global model’s ability to learn from the data of the remaining clients.
IV-B Collaborative Dictator Clients
We next examine the impact of coordinated attacks involving multiple dictator clients. In this setting, two or three clients jointly follow the attack strategy, described in Algorithm 2, aiming to eliminate the influence of all other participants. Table I and Figure 2 summarize the outcomes. The results demonstrate that the collaborating dictator clients succeed in entirely erasing the influence of the benign clients, leading the global model to achieve $0.00\%$ accuracy on their data. Simultaneously, the global model maintains high accuracy on the data held by the collaborative dictators, indicating that it has effectively converged to a model tailored solely to their objectives. These findings further reinforce the practicality and scalability of our proposed attack strategy in multi-attacker scenarios. The coordinated behavior among the dictator clients allows them to dominate the training process, ensuring that the global model exclusively reflects their data distributions while ignoring the contributions of the remaining benign participants. The success of this attack highlights the vulnerability of FL even when malicious clients are in minority, provided they act in collaboration.
V Competition and Collusion Among Dictator Clients
We also explore the nuanced interactions that can arise among dictator clients in FL systems. In Appendix F, we examine a competitive setting where every participating client independently aims to become the sole dictator and dominate the global model—an extreme yet insightful scenario. Furthermore, in Appendix G, we explore a more collaborative dynamic, where multiple dictator clients form alliances. We investigate whether such cooperation is inherently stable or if, ultimately, some clients can strategically betray their collaborators to gain a greater influence over the model.
Finally, to account for more realistic training scenarios, we conduct experiments in which a randomly selected client is dropped from each update round (Appendix L). We also evaluate the robustness of our attack against server-side gradient norm clipping [18] as a defense (Appendix M).
VI Conclusion
Appendix A Figures
Figure 5: Loss functions for mutual domination scenario
Figure 6: Loss functions for betrayal in collaboration scenario
Appendix B Algorithm 1: Single Dictator Client
@@ Appendix C Algorithm 2: Collaborative Dictator Clients
Appendix D Algorithm 3: Cheater Client
@@
Appendix E Practical Implications
Appendix F Mutual Domination: When Every Client Seeks Control
Here, we explore the scenario where all clients independently act as dictators, each attempting to retain only its own contribution while nullifying the influence of others. In other words, each client follows the update strategy outlined in Algorithm 1. In practice, such behavior leads to a catastrophic failure of learning, with the global model failing to converge and the loss growing exponentially. We analyze the underlying reason behind this phenomenon in what follows. At iteration 0, the server sends the initialized weights $\theta_{0}$ to all clients. Each client then computes its local gradient, and the server aggregates these to update the global model as $\theta_{1}=\theta_{0}-\eta\sum\nolimits_{n=1}^{N}\nabla\mathcal{L}_{n}(\theta_{0})$. In the next iteration, the server broadcasts $\theta_{1}$ to all clients. Now, each client attempts to simulate what the global model would have been if it alone had contributed to the update. For each client $n\in\mathcal{N}$, we define $\hat{\theta}^{n}_{1}$ as the hypothetical global model after iteration 0 only if client $n$ had participated. Using this, each client computes its malicious update $M^{n}_{1}$, as defined in Section III-A as $M^{n}_{1}=\nabla{\mathcal{L}_{n}}(\hat{\theta}^{n}_{1})-\left(\frac{\theta_{0}-\theta_{1}}{\eta}-\nabla\mathcal{L}_{n}(\theta_{0})\right)$. Now, we analyze the updated global model $\theta_{2}$ after the server aggregates the updates from all clients in the second iteration:
$\theta_{2}$ $=\theta_{1}-\eta\sum\nolimits_{n=1}^{N}M^{n}_{1}=\theta_{1}-\eta(\sum\nolimits_{n=1}^{N}\nabla{\mathcal{L}_{n}}(\hat{\theta}^{n}_{1})$ (7)
$\quad-(\frac{\theta_{0}-\theta_{1}}{\eta}-\nabla\mathcal{L}_{n}(\theta_{0})))$
$=\theta_{1}-\eta(\sum\nolimits_{n=1}^{N}\nabla{\mathcal{L}_{n}}(\hat{\theta}^{n}_{1})-\frac{N(\theta_{0}-\theta_{1})}{\eta}$ (8)
$\quad+\sum\nolimits_{n=1}^{N}\nabla{\mathcal{L}_{n}(\theta_{0})})$
$=\theta_{0}-\eta\sum\nolimits_{n=1}^{N}\nabla{\mathcal{L}_{n}(\theta_{0})}+\eta(N-1)\sum\nolimits_{n=1}^{N}\nabla{\mathcal{L}_{n}(\theta_{0})}$ (10)
$\quad-\eta\sum\nolimits_{n=1}^{N}\nabla{\mathcal{L}_{n}(\hat{\theta}^{n}_{1}})$
Since N2>0N-2>0 assuming that we have more than $2$ clients in the system, and the learning rate $\eta$ is a positive real number, it follows that $\eta(N-2)>0$. Consequently, from Eq. (11), it follows that when all clients act as independent dictators and send the defined malicious update, the resulting model update effectively moves in the opposite direction of intended gradient. In other words, the updating procedure resembles gradient ascent instead of gradient descent, and thereby increasing the loss rather than minimizing it. This behavior causes the model to ”unlearn” the progress made in previous iteration. Therefore, when every client behaves as an independent dictator, the global model fails to learn meaningful representations and make no effective progress. it unlearns the knowledge acquired in the previous iteration. Therefore, in the scenario where every client is an independent dictator, the global model will learn almost nothing. Our empirical results, presented in Section F-A, confirms this breakdown in learning in practice.
F-A Experiments for Mutual Domination
We now consider the extreme scenario where every client behaves as an independent dictator, each executing Algorithm 1 to preserve only its own contribution while nullifying the effects of all others. As established theoretically in Section F, this adversarial configuration results in mutually destructive behavior, where no single client’s update can effectively influence the global model without being canceled out by others, resulting in a destructive equilibrium where no useful learning can occur. The empirical results, shown in Figure 6 (in Appendix A), strongly corroborate this. The global model fails to make progress on any client’s data; instead of converging, the loss increases rapidly and consistently across all datasets. This behavior aligns with the theoretical finding that the aggregated updates approximate a form of gradient ascent, undoing prior learning and leading to model divergence. This experiment underscores a key insight: when all clients prioritize their own influence at the expense of others, the entire system collapses. FL becomes ineffective, highlighting the need for defenses against not only isolated attackers but also adversarial groups.
Appendix G Betrayal in Collaboration: Strategic Cheating Among Dictator Clients
Here, we show that even collaborative dictators—those collaborating to erase other participants’ contributions—may ultimately betray one another. For simplicity, we focus on a setting where the set of collaborative dictator clients is $\mathcal{P}=\{1,2\}$. As illustrated in Figure 4 (in Appendix A), we consider the case where dictator client $1$, decides to cheat its partner, client $2$, after a specific iteration $E$. While both clients initially cooperate using Algorithm 2 to jointly eliminate the influence of all other clients, we introduce Algorithm 3 (in Appendix D), which enables client $1$ to unilaterally eliminate client $2$’s contribution as well, effectively taking full control of the model.
Prior to iteration $E$, client $1$ shares its gradients with client $2$, contributing jointly to a local model $\hat{\theta}^{\mathcal{P}}_{t}$. However, simultaneously, client $1$ secretly maintains a private model $\hat{\theta}^{1}_{t}$, which simulates the evolution of the global model if only client $1$ participated in training. At each iteration, client $1$ computes a correction term $\Delta_{t}=\nabla{\mathcal{L}_{1}(\hat{\theta}^{m}_{t})}-(\nabla{\mathcal{L}_{1}(\hat{\theta}^{\mathcal{P}}_{t})}+\nabla{\mathcal{L}_{2}(\hat{\theta}^{\mathcal{P}}_{t})})$, which captures the discrepancy between acting alone and collaborating. These differences are accumulated into a cheating offset, denoted as $\mathrm{Cheating\_Update}$. At iteration $E$, client $1$ sends this accumulated update to the server in place of the expected collaborative update. This forces the server to jump to a state equivalent to one where if only client $1$ had participated throughout the training process—effectively eliminating the contribution of client $2$, despite their prior collaboration, as well as the benign clients’ influence. A formal proof of this strategy is provided in Appendix J; our empirical results in Section G-A confirm the effectiveness of this betrayal strategy in practice.
G-A Experiments for Betrayal in Collaboration
In this experiment, we evaluate a scenario in which two clients, client $1$ and client $2$, initially act as collaborative dictators. While both begin by coordinating via Algorithm 2, client $1$ eventually deviates and follows the betrayal strategy outlined in Algorithm 3 (discussed in Section G). This setup allows client $1$ to secretly prepare for a unilateral takeover of the model. As shown in Figure 6 (in Appendix A), at iteration 10—the predetermined betrayal point—the global model abruptly loses performance on client $2$’s dataset, while having even lower loss on client $1$’s data. This result confirms that client $1$ successfully erases not only the contributions of the benign clients, but also those of its former collaborator, client $2$. These findings empirically validate that a malicious client can strategically cooperate to gain trust, only to later betray its partners and assert full control over the global model. This highlights a critical vulnerability in FL; even collaborative adversaries can be exploited by more sophisticated attackers acting within their own group.